Eighteen years ago a woman was 1/2 as old as her husband, while 4 years ago she was 2/3 as old as her husband. Find both of their ages now.
This is how far I have gotten:
w=wife's current age
h= husbands current age
(w-18)=(h-18)/2
Not even sure if I'm on the right track.....

Answer by josgarithmetic(39617)

w-18=(1/2)(h-18), for Eighteen years ago a woman was 1/2 as old as her husband,
w-4=(2/3)(h-4), for 4 years ago she was 2/3 as old as her husband.

You WERE on the right track.